In a recent year, the total sales at restaurant A were $450,000 more than at restaurant B. Total sales at the two stores were $600,000,000. What were the sales for the year at each store?
1
Expert's answer
2013-03-19T09:22:35-0400
Suppose, the total sales at restaurant A were $X. Then, the total sales at restaurant B were $(X-450,000).
Total sales at the two stores were $600,000,000:
X + (X-450,000) = 600,000,000
Solving this equation:
2X = 600,000,000 + 450,000
X = 600,450,000/2 = 300225000
Therefore, the total sales at restaurant A were $300,225,000.
For restaurant B we have:
300225000 - 450,000 = 299775000
Therefore, the total sales at restaurant B were $299,775,000.
Answer: the total sales at restaurant A were $300,225,000, at restaurant B - $299,775,000.